# Benchmarking Universal Machine Learning Interatomic Potentials for Elastic Property Prediction

### v2

@@ Abstract
Universal machine learning interatomic potentials have emerged as efficient tools for materials simulation, yet their reliability for elastic property prediction remains unclear. Here, we present a systematic benchmark of four uMLIPs—MatterSim, MACE, SevenNet, and CHGNet—against first-principles data for nearly 11 000 elastically stable materials from the Materials Project database. The results show that SevenNet achieves the highest accuracy, MACE and MatterSim balance accuracy with efficiency, while CHGNet performs less effectively overall. To further improve predictive quality, we perform targeted fine-tuning on all four uMLIPs using strained configurations derived from 185 high-error materials. After fine-tuning, CHGNet exhibits the largest overall improvement, with an average mean absolute percentage error reduction of about 23%, followed by MatterSim at around 21% and SevenNet at 18%, whereas MACE shows a performance degradation of roughly 14%. This work provides quantitative guidance for model selection and data refinement, advancing uMLIPs toward reliable applications in mechanical property prediction.
I Introduction
@@
Elastic properties [10.1115/1.3423687], as fundamental properties of materials, play an important role in governing their mechanical behavior across a wide range of applications, from structural engineering to lithium battery systems [crystal3, Kim_2021, XIE2021104602], and related fields [phonons]. Accurate prediction of elastic constants and their derived mechanical parameters, such as bulk modulus, shear modulus, Young’s modulus, and Poisson’s ratio, is a critical task of computational materials design [YU2010671]. Although the modern density functional theory (DFT) [hafner2006toward, hohenberg1964inhomogeneous, kohn1965self, curtarolo2005accuracy] provides reliable and reproducible predictions of elastic properties, such calculations are often associated with high computational costs in high-throughput materials screening. Owing to this computational bottleneck, the systematic exploration of large chemical spaces is strictly constrained [de2015charting], which in turn hinders the efficient evaluation of elastic mechanical properties and delays materials design and discovery.
@@
In recent years, machine learning interatomic potentials (MLIPs) [mueller2020machine, friederich2021machine, gokcan2022learning, deringer2019machine, botu2017machine] have rapidly emerged as important tools in materials simulation, offering an effective balance between the high accuracy of quantum mechanical calculations and the efficiency of classical potentials. Generally, these models are obtained by learning interatomic interactions from large-scale DFT data sets, and enable predictions with near-quantum accuracy while substantially reducing computational cost for crystal structure prediction [crystal1, crystal2, crystal3], molecular dynamics simulation [MD1, MD2], and related tasks [related1, related2, cost]. Recent advances in graph neural networks, message-passing architectures, and equivariant representations have greatly improved the capabilities of MLIPs. These developments have led to the emergence of universal MLIPs (uMLIPs) [doi:10.1021/acsami.4c03815, chen2019graph, chen2022universal, deng2023chgnet], which can accurately model a wide range of chemical compositions and crystal structures. However, accurately predicting elastic properties requires a dependable evaluation of the second derivatives of the potential energy surface (PES), which introduces stricter and qualitatively different challenges than those encountered in predicting energies and forces.
@@
So far, many efforts have been devoted to developing uMLIPs that improve the accuracy of energy, force, and stress predictions [deng2025systematic]. For instance, previous studies have shown that uMLIPs on the Matbench platform [dunn2020benchmarking] perform well in structural optimization, structure prediction, and molecular dynamics simulations. However, their reliability and effectiveness in predicting elastic properties remain unexplored. This is because the relationship between energy–force accuracy and second-derivative precision is not straightforward, as elastic constants are highly sensitive to slight variations in the curvature of the PES, which are often difficult to capture with conventional training strategy. Therefore, analyzing the difference between the overall predictive accuracy and property-specific performance of uMLIPs, and evaluating their capability in mechanical property predictions, is of great importance.
@@
In this work, we conduct a systematic evaluation to address the existing gap in crystal mechanical property research. Specifically, we employ four universal machine learning interatomic potentials (uMLIPs) — Crystal Hamiltonian Graph Neural Network (CHGNet) [deng2023chgnet], MACE [batatia2022mace], MatterSim [yang2024mattersim], and Scalable EquiVariance-Enabled Neural Network (SevenNet) [park2024scalable] — to calculate the elastic properties of 10 994 crystal structures from the Materials Project database [horton2025accelerated, jain2013commentary, de2015charting, jain2011formation, wang2021framework, aykol2018thermodynamic, ong2008li], and systematically compare the results with the DFT reference data provided therein. We further quantify model performance differences in key indicators such as shear modulus, bulk modulus, Young’s modulus, Poisson’s ratio, and mechanical stability, as well as computational efficiency. Moreover, we introduce a targeted fine-tuning scheme that augments uMLIP training with strained configurations from high-error materials, enabling a direct evaluation of how incorporating non-equilibrium data affects mechanical predictive accuracy. Building on these analyses, we propose evidence-based guidelines for the suitable selection of uMLIPs and demonstrate that targeted fine-tuning can further enhance their accuracy and reliability in predicting elastic properties.
II Methodology
II.1 Dataset Construction and Analysis
@@
Figure 1: Crystal structure analysis of the dataset. (a) Periodic table heatmap indicating element occurrence. (b) Sunburst plot illustrating the distribution of crystal systems and space groups, with integers placed at the outer margins representing the corresponding space-group numbers. (c) Histogram of atom counts per unit cell.
@@
In this work, we collected 10,994 structures with reported elastic properties from the Materials Project database. Among them, 10,871 structures were mechanically stable at the DFT level, and these were used as our benchmark dataset. In Fig. 1(a), we present the distribution of elements. The nonmetals such as B, C, N, and O, main-group metals like Li and Mg, and transition metals including Ni, Cu, Zn, and Ti appear most often. Heavy and radioactive elements are rare. From a crystallographic perspective (see Fig. 1(b)), the dataset covers seven crystal systems. Cubic structures are the most common (23%), followed by tetragonal (20%) and orthorhombic (19%). Trigonal and monoclinic systems make up 16% and 12%, while hexagonal (7%) and triclinic (3%) systems are less frequent. In total, 169 space groups are represented, giving wide crystallographic diversity. Finally, from the distribution of the number of atoms in Fig. 1(c), we find that most structures have fewer than 20 atoms per unit cell, with 5-10 atoms being the most typical, and structures with more than 30 atoms are uncommon.
@@
In Fig. 2, we present the distribution of the band gap, formation energy, and basic mechanical properties of the dataset, showing that these quantities exhibit a broad and diverse distribution. The statistical analysis shows that 3,248 materials (29.9%) are semiconductors or insulators, with an average band gap of 0.69 eV, while the remaining 7,623 materials (70.1%) are metallic. For the semiconductor subset, as illustrated in Fig. 2(a), the majority of structures possess negative formation energies (mean: -0.90 ± 0.98 eV/atom) and energy above hull values (mean: 0.03 ± 0.10 eV/atom) close to zero, indicating good thermodynamic stability. Regarding mechanical properties in Fig. 2(b), the dataset shows that the bulk moduli range from 0.33 to 491.33 GPa (mean: 104.41 ± 73.73 GPa), shear moduli from 0.45 to 525.42 GPa (mean: 50.93 ± 44.22 GPa), and Poisson’s ratios from -0.48 to 0.80 (mean: 0.29 ± 0.07). Overall, the dataset demonstrates strong representativeness in electronic, thermodynamic, and mechanical domains, providing a reliable sample for evaluating elastic properties in real materials.
@@
Figure 2: Scatter plot of formation energy versus band gap, color-coded by the energy above hull. Marginal histograms illustrate the distribution of formation energies and band gaps. (b) Elastic property correlations. Scatter plot of bulk modulus versus shear modulus (log scale), color-coded by Poisson’s ratio. Marginal histograms show the distributions of bulk and shear moduli.
II.2 Brief Description of Evaluated uMLIPs
@@
In CHGNet [deng2023chgnet], the total potential energy is expressed as
$E_{\text{tot}}=\sum_{i}\;L_{3}\circ g\circ L_{2}\circ g\circ L_{1}\big(\mathbf{v}_{i}^{(4)}\big)$ (1)
@@
where $L_{1}$, $L_{2}$, and $L_{3}$ are successive linear transformations, and $g$ is a nonlinear activation function (typically the SiLU function). The vector $\mathbf{v}_{i}^{(4)}$ represents the final latent feature of atom $i$, obtained after four message-passing layers that aggregate both local bonding environments and longer-range structural correlations. Through this hierarchical transformation, each atomic environment is mapped to a high-dimensional representation that captures the coupling between geometric and electronic degrees of freedom. The total energy $E_{\text{tot}}$ is then constructed as a smooth, differentiable function of all atomic positions, ensuring physical consistency between predicted energies, forces, and stresses. CHGNet enhances this framework by embedding charge information into the latent space via magnetic moment constraints, which effectively incorporate electronic-structure effects into the learned potential. This charge-informed representation enables the model to distinguish between different ionic states and electronic configurations, a capability essential for accurately describing materials where charge redistribution and orbital occupancy govern structural stability, phase behavior, and transport properties.
@@
MACE [batatia2022mace] advances interatomic potential modeling by combining the systematic completeness of Atomic Cluster Expansion (ACE) with the higher-order equivariant message passing of modern graph neural networks. Unlike conventional message-passing neural networks that primarily encode two-body interactions and rely on deep stacking to capture higher-order correlations, MACE constructs explicit many-body messages within each layer through a hierarchical expansion,
$\begin{split}m_{i}^{(t)}=&\sum_{j}u_{1}\!\big(\sigma_{i}^{(t)};\sigma_{j}^{(t)}\big)+\sum_{j_{1},j_{2}}u_{2}\!\big(\sigma_{i}^{(t)};\sigma_{j_{1}}^{(t)},\sigma_{j_{2}}^{(t)}\big)+\cdots\\ &+\sum_{j_{1},\dots,j_{\nu}}u_{\nu}\!\big(\sigma_{i}^{(t)};\sigma_{j_{1}}^{(t)},\dots,\sigma_{j_{\nu}}^{(t)}\big)\end{split}$ (2)
where $m_{i}^{(t)}$ is the message received by atom $i$ at layer $t$, $\sigma_{i}^{(t)}=(\mathbf{r}_{i},z_{i},h_{i}^{(t)})$ denotes its geometric, chemical, and latent state, and $u_{\nu}$ are learnable tensorial functions encoding correlations up to body order $(\nu+1)$. This formulation embeds the full hierarchy of local many-body interactions directly into each message-passing step, making the representation both equivariant under $E(3)$ transformations and systematically improvable by increasing the correlation order $\nu$. As a result, MACE achieves the accuracy of high-order ACE potentials with only two network layers, while maintaining linear scaling, GPU-friendly parallelism, and full physical symmetry. This fusion of traditional many-body theory and modern equivariant learning enables MACE to deliver quantum-level precision and computational efficiency, bridging the gap between explicit many-body potentials and scalable neural force-field architectures.
@@
The MatterSim potential [yang2024mattersim] is a large-scale, symmetry-preserving machine-learning force field that combines the M3GNet architecture with a periodic-aware Graphormer backbone. Each atomic structure is represented as a graph $G=(Z,V,R,[L,S])$, where atomic nodes $Z$ carry feature vectors $v_{i}$, edges $V$ connect atom pairs $(i,j)$ within a cutoff radius $r_{c}$, $R=\{r_{i}\}$ are atomic coordinates, and $[L,S]$ encodes the global lattice and thermodynamic state. In the M3GNet message-passing block, each edge feature $e_{ij}$ represents the bond between atoms $i$ and $j$, including pairwise information such as chemical type and interatomic distance $r_{ij}=\|r_{i}-r_{j}\|$. To incorporate three-body geometry, $e_{ij}$ is refined through a spherical-Bessel / spherical-harmonic expansion of its neighboring environment:
$\tilde{e}_{ij}=\sum_{k}j_{\ell}\!\Big(\tfrac{z_{\ell n}\|r_{ik}\|}{r_{c}}\Big)Y_{\ell}^{0}(\theta_{jik})\otimes\sigma(W_{v}v_{k}+b_{v})f_{c}(\|r_{ij}\|)f_{c}(\|r_{ik}\|)$ (3)
@@
where $r_{ij}=r_{i}-r_{j}$ and $\theta_{jik}$ denotes the angle between bonds $e_{ij}$ and $e_{ik}$; the functions $j_{\ell}$ and $Y_{\ell}^{0}$ correspond to spherical Bessel functions and spherical harmonics with roots $z_{\ell n}$,
$f_{c}(r)=1-6(r/r_{c})^{5}+15(r/r_{c})^{4}-10(r/r_{c})^{3}$
is a smooth cutoff ensuring continuity at $r_{c}$, and $W,b$ are learnable weights and biases. The intermediate term $\tilde{e}_{ij}$ aggregates angular information from neighboring atoms $k$, and the updated edge feature $e^{\prime}_{ij}$ is obtained through nonlinear mixing:
$e^{\prime}_{ij}=e_{ij}+g(\tilde{W}_{2}\tilde{e}_{ij}+\tilde{b}_{2})\otimes\sigma(\tilde{W}_{1}\tilde{e}_{ij}+\tilde{b}_{1})$ (4)
where $\sigma$ is the sigmoid activation and $g(x)=x\,\sigma(x)$. Here, $e_{ij}$ encodes pairwise interactions, $\tilde{e}_{ij}$ introduces three-body angular geometry, and $e^{\prime}_{ij}$ forms the refined many-body bond embedding passed to the next layer. Built upon these physically grounded descriptors and extended with long-range, periodic-aware attention, MatterSim achieves robust generalization and order-of-magnitude accuracy improvements over previous universal machine-learning force fields, trained on more than 17 million first-principles structures spanning diverse compositions and thermodynamic conditions.
@@
SevenNet [park2024scalable] (Scalable EquiVariance-Enabled Neural NETwork) follows the atom-decomposed energy formalism widely used in machine-learned interatomic potentials. This locality ensures that the computational cost scales linearly with the number of atoms, $\mathcal{O}(N)$, enabling large-scale molecular dynamics with thousands to millions of atoms. At each message-passing layer $t$, atomic features are updated by
@@
where $M_{t}$ and $U_{t}$ are learnable equivariant mappings that propagate geometric information between atoms while preserving rotational and permutational symmetry. The edge feature ${e}_{vw}^{(t)}$ is constructed from the relative displacement vector ${r}_{vw}={r}_{w}-{r}_{v}$ and encodes both its magnitude and orientation, ensuring proper transformation under three-dimensional rotations. SevenNet extends the NequIP architecture by reorganizing its forward and reverse communication for efficient spatial domain decomposition.
II.3 Elastic Property Calculations
The second-order elastic constants ($C_{ij}$) were calculated using the stress-strain method [TBarron_1965]. According to Hooke’s law, the relationship between stress $\sigma_{ij}$ and strain $\varepsilon_{kl}$ with Voigt notation can be expressed as:
$\sigma_{i}=C_{ij}\varepsilon_{j}\quad(i,j=\{1,2,3,4,5,6\})$ (7)

### v3

@@ Abstract
Universal machine learning interatomic potentials (uMLIPs) have emerged as efficient tools for materials simulation, yet their reliability for elastic property prediction remains unclear. Here, we present a systematic benchmark of four uMLIPs—MatterSim, MACE, SevenNet, and CHGNet—against first-principles data for nearly 11 00011\,000 elastically stable materials from the Materials Project database. The results show that SevenNet achieves the highest accuracy, MACE and MatterSim balance accuracy with efficiency, while CHGNet performs less effectively overall. To further improve predictive quality, we perform targeted fine-tuning on all four uMLIPs using strained configurations derived from $185$ high-error materials. After fine-tuning, CHGNet shows the most substantial improvement in overall accuracy, with MatterSim and SevenNet also benefiting from the fine-tuning, whereas MACE shows limited robustness to this procedure. This work provides quantitative guidance for model selection and data refinement, advancing uMLIPs toward reliable applications in mechanical property prediction.
keywords:
Universal machine learning interatomic potentials , Elastic properties , Materials Project , Fine-tuning , Mechanical stability
journal: Mechanics of Materials
[…]
[inst1]organization=School of Intelligence Science and Technology, Nanjing University of Science and Technology, addressline=, city=Jiangyin, state=Jiangsu, postcode=214443, country=China
[…]
[inst2]organization=School of Physics, Hefei University of Technology, addressline=, city=Hefei, state=Anhui, postcode=230009, country=China
1 Introduction
@@
Elastic properties [1], as fundamental properties of materials, play an important role in governing their mechanical behavior across a wide range of applications, from structural engineering to lithium battery systems [2, 3, 4], and related fields [5]. Accurate prediction of elastic constants and their derived mechanical parameters, such as bulk modulus, shear modulus, Young’s modulus, and Poisson’s ratio, is a critical task of computational materials design [6]. Although modern density functional theory (DFT) [7, 8, 9, 10] provides reliable and reproducible predictions of elastic properties, such calculations are often associated with high computational costs in high-throughput materials screening. Owing to this computational bottleneck, the systematic exploration of large chemical spaces is strictly constrained [11], which in turn hinders the efficient evaluation of elastic mechanical properties and delays materials design and discovery.
@@
In recent years, machine learning interatomic potentials (MLIPs) [12, 13, 14, 15, 16] have rapidly emerged as important tools in materials simulation, offering an effective balance between the high accuracy of quantum mechanical calculations and the efficiency of classical potentials. Generally, these models are obtained by learning interatomic interactions from large-scale DFT data sets, and enable predictions with near-quantum accuracy while substantially reducing computational cost for crystal structure prediction [17, 18, 2], molecular dynamics simulation [19, 20], and related tasks [21, 22, 23]. Recent advances in graph neural networks, message-passing architectures, and equivariant representations have greatly improved the capabilities of MLIPs. These developments have led to the emergence of universal machine learning interatomic potentials (uMLIPs) [24, 25, 26, 27], which can accurately model a wide range of chemical compositions and crystal structures. However, accurately predicting elastic properties requires a dependable evaluation of the second derivatives of the potential energy surface (PES), which introduces stricter and qualitatively different challenges than those encountered in predicting energies and forces.
@@
So far, many efforts have been devoted to developing uMLIPs that improve the accuracy of energy, force, and stress predictions [28]. For instance, previous studies have shown that uMLIPs on the Matbench platform [29] perform well in structural optimization, structure prediction, and molecular dynamics simulations. However, their reliability and effectiveness in predicting elastic properties remain unexplored. This is because the relationship between energy–force accuracy and second-derivative precision is not straightforward, as elastic constants are highly sensitive to slight variations in the curvature of the PES, which are often difficult to capture with conventional training strategy. Therefore, analyzing the difference between the overall predictive accuracy and property-specific performance of uMLIPs, and evaluating their capability in mechanical property predictions, is of great importance.
@@
In this work, we conduct a systematic evaluation to address the existing gap in crystal mechanical property research. Specifically, we employ four uMLIPs — Crystal Hamiltonian Graph Neural Network (CHGNet) [27], MACE [30], MatterSim [31], and Scalable EquiVariance-Enabled Neural Network (SevenNet) [32] — to calculate the elastic properties of 10 99410\,994 crystal structures from the Materials Project database [33, 34, 11, 35, 36, 37, 38], and systematically compare the results with the DFT reference data provided therein. We further quantify model performance differences in key indicators such as shear modulus, bulk modulus, Young’s modulus, Poisson’s ratio, and mechanical stability, as well as computational efficiency. Moreover, we introduce a targeted fine-tuning scheme that augments uMLIP training with strained configurations from high-error materials, enabling a direct evaluation of how incorporating non-equilibrium data affects mechanical predictive accuracy. Building on these analyses, we propose evidence-based guidelines for the suitable selection of uMLIPs and demonstrate that targeted fine-tuning can further enhance their accuracy and reliability in predicting elastic properties.
2 Methodology
2.1 Dataset Construction and Analysis
@@
Figure 1: Crystal structure analysis of the dataset. (a) Periodic table heatmap indicating element occurrence. (b) The distribution of crystal systems and space groups, with integers placed at the outer margins representing the corresponding space-group numbers. (c) Histogram of atom counts per unit cell.
@@
In this work, we collected 10 99410\,994 structures with reported elastic properties from the Materials Project database. Among them, 10 87110\,871 structures were mechanically stable at the DFT level, and these were used as our benchmark dataset. In Fig. 1(a), we present the distribution of elements. The nonmetals such as B, C, N, and O, main-group metals like Li and Mg, and transition metals including Ni, Cu, Zn, and Ti appear most often. Heavy and radioactive elements are rare. From a crystallographic perspective (see Fig. 1(b)), the dataset covers seven crystal systems. Cubic structures are the most common (23%), followed by tetragonal (20%) and orthorhombic (19%). Trigonal and monoclinic systems make up 16% and 12%, while hexagonal (7%) and triclinic (3%) systems are less frequent. In total, 169 space groups are represented, giving wide crystallographic diversity. Finally, from the distribution of the number of atoms in Fig. 1(c), we find that most structures have fewer than 20 atoms per unit cell, with 5–10 atoms being the most typical, and structures with more than 30 atoms are uncommon.
@@
In Fig. 2, we present the distribution of the band gap, formation energy, and basic mechanical properties of the dataset, showing that these quantities exhibit a broad and diverse distribution. The statistical analysis shows that 3 2483\,248 materials (29.9%) are semiconductors or insulators, with an average band gap of 0.69 eV, while the remaining 7 6237\,623 materials (70.1%) are metallic. For the semiconductor subset, as illustrated in Fig. 2(a), the majority of structures possess negative formation energies (mean: $0.90\pm 0.98$ eV/atom) and energy above hull values (mean: $0.03\pm 0.10$ eV/atom) close to zero, indicating good thermodynamic stability. Regarding mechanical properties in Fig. 2(b), the dataset shows that the bulk moduli range from 0.33 to 491.33 GPa (mean: $104.41\pm 73.73$ GPa), shear moduli from 0.45 to 525.42 GPa (mean: $50.93\pm 44.22$ GPa), and Poisson’s ratios from 0.48-0.48 to 0.80 (mean: $0.29\pm 0.07$). Overall, the dataset demonstrates strong representativeness in electronic, thermodynamic, and mechanical domains, providing a reliable sample for evaluating elastic properties in real materials.
@@
Figure 2: (a) Scatter plot of formation energy versus band gap, color-coded by the energy above hull. Marginal histograms illustrate the distribution of formation energies and band gaps. (b) Elastic property correlations. Scatter plot of bulk modulus versus shear modulus (log scale), color-coded by Poisson’s ratio. Marginal histograms show the distributions of bulk and shear moduli.
2.2 Brief Description of Evaluated uMLIPs
@@
In CHGNet [27], the total potential energy is expressed as
$E_{\text{tot}}=\sum_{i}\;L_{3}\circ g\circ L_{2}\circ g\circ L_{1}\big(\mathbf{v}_{i}^{(4)}\big),$ (1)
@@
where $L_{1}$, $L_{2}$, and $L_{3}$ are successive linear transformations, and $g$ is a nonlinear activation function (typically the SiLU function). The vector $\mathbf{v}_{i}^{(4)}$ represents the final latent feature of atom $i$, obtained after four message-passing layers that aggregate both local bonding environments and longer-range structural correlations. Through this hierarchical transformation, each atomic environment is mapped to a high-dimensional representation that captures the coupling between geometric and electronic degrees of freedom. The total energy $E_{\text{tot}}$ is then constructed as a smooth, differentiable function of all atomic positions, ensuring physical consistency between predicted energies, forces, and stresses. CHGNet enhances this framework by embedding charge information into the latent space via magnetic moment constraints, which effectively incorporate electronic-structure effects into the learned potential.
@@
MACE [30] advances interatomic potential modeling by combining the systematic completeness of Atomic Cluster Expansion (ACE) with the higher-order equivariant message passing of modern graph neural networks. Unlike conventional message-passing neural networks that primarily encode two-body interactions and rely on deep stacking to capture higher-order correlations, MACE constructs explicit many-body messages within each layer through a hierarchical expansion,
$\begin{split}m_{i}^{(t)}=&\sum_{j}u_{1}\!\big(\sigma_{i}^{(t)};\sigma_{j}^{(t)}\big)+\sum_{j_{1},j_{2}}u_{2}\!\big(\sigma_{i}^{(t)};\sigma_{j_{1}}^{(t)},\sigma_{j_{2}}^{(t)}\big)+\cdots\\ &+\sum_{j_{1},\dots,j_{\nu}}u_{\nu}\!\big(\sigma_{i}^{(t)};\sigma_{j_{1}}^{(t)},\dots,\sigma_{j_{\nu}}^{(t)}\big),\end{split}$ (2)
where $m_{i}^{(t)}$ is the message received by atom $i$ at layer $t$, $\sigma_{i}^{(t)}=(\mathbf{r}_{i},z_{i},h_{i}^{(t)})$ denotes its geometric, chemical, and latent state, and $u_{\nu}$ are learnable tensorial functions encoding correlations up to body order $(\nu+1)$.
@@
The MatterSim potential [31] is a large-scale, symmetry-preserving machine-learning force field that combines the M3GNet architecture with a periodic-aware Graphormer backbone. Each atomic structure is represented as a graph $G=(Z,V,R,[L,S])$, where atomic nodes $Z$ carry feature vectors $v_{i}$, edges $V$ connect atom pairs $(i,j)$ within a cutoff radius $r_{c}$, $R=\{r_{i}\}$ are atomic coordinates, and $[L,S]$ encodes the global lattice and thermodynamic state. In the M3GNet message-passing block, each edge feature $e_{ij}$ represents the bond between atoms $i$ and $j$, including pairwise information such as chemical type and interatomic distance $r_{ij}=\|r_{i}-r_{j}\|$. To incorporate three-body geometry, $e_{ij}$ is refined through a spherical-Bessel / spherical-harmonic expansion of its neighboring environment:
$\tilde{e}_{ij}=\sum_{k}$ $j_{\ell}\!\Big(\tfrac{z_{\ell n}\lVert r_{ik}\rVert}{r_{c}}\Big)\,Y_{\ell}^{0}(\theta_{jik})\otimes\sigma(W_{v}v_{k}+b_{v})$ (3)
$\qquad\times f_{c}(\lVert r_{ij}\rVert)\,f_{c}(\lVert r_{ik}\rVert)\,.$
@@
where $\theta_{jik}$ denotes the angle between bonds $e_{ij}$ and $e_{ik}$. The smooth cutoff is
$f_{c}(r)=1-6(r/r_{c})^{5}+15(r/r_{c})^{4}-10(r/r_{c})^{3},$ (4)
and the updated edge feature is obtained through nonlinear mixing:
$e^{\prime}_{ij}=e_{ij}+g(\tilde{W}_{2}\tilde{e}_{ij}+\tilde{b}_{2})\otimes\sigma(\tilde{W}_{1}\tilde{e}_{ij}+\tilde{b}_{1}),$ (5)
where $\sigma$ is the sigmoid activation and $g(x)=x\,\sigma(x)$.
@@
SevenNet [32] follows the atom-decomposed energy formalism widely used in machine-learned interatomic potentials. This locality ensures that the computational cost scales linearly with the number of atoms, $\mathcal{O}(N)$, enabling large-scale molecular dynamics. At each message-passing layer $t$, atomic features are updated according to
@@
Here, $M_{t}$ and $U_{t}$ are learnable equivariant mappings that propagate geometric information between atoms while preserving rotational and permutational symmetries.
2.3 Elastic Property Calculations
The second-order elastic constants ($C_{ij}$) were calculated using the stress–strain method [39]. According to Hooke’s law, the relationship between stress $\sigma_{ij}$ and strain $\varepsilon_{kl}$ with Voigt notation can be expressed as:
$\sigma_{i}=C_{ij}\varepsilon_{j}\quad(i,j=\{1,2,3,4,5,6\}).$ (8)